OpenStudy (anonymous):

Solve 3/(X^2-7x+10) + 2 = (x-4)/ (x-5) and eplain. thanksss!

OpenStudy (dumbcow):

first Factor: factors of 10 that add up to -7 \[x^{2}-7x+10 = (x-5)(x-2)\] add fractions getting common denominator \[\frac{3}{(x-5)(x-2)} + \frac{2(x-5)(x-2)}{(x-5)(x-2)} = \frac{3+2(x-5)(x-2)}{(x-5)(x-2)}\] expand numerator by distributing and adding like terms \[=\frac{2x ^{2}-14x+23}{(x-5)(x-2)}\] Notice 2 fractions equal to each other, we can cross-multiply \[(x-5)(2x^{2}-14x+23) = (x-4)(x-5)(x-2)\] Divide by (x-5) on both sides, they cancel FOIL right side, \[2x^{2}-14x+23 = x^{2}-6x+8\] move everything to left side so it equals 0 \[x^{2} -8x +15 =0\] Factor and solve \[(x-5)(x-3) = 0\] \[x=5, x=3\]